# Guiding diffusion models to reconstruct flow fields from sparse data

### v1

@@ IV Conclusions
We have shown that our method outperforms previous diffusion-based reconstruction approaches in both 2D and 3D scenarios. In particular, it excels at producing physically consistent reconstructions while preserving key flow statistics and accurately capturing the smallest scales.
Our results point to several interesting avenues for future work, such as evaluating how the proposed generative method performs for reconstructing other flow problems. Additionally, the generalization capabilities of our trained models to different Reynolds numbers and boundary conditions, as well as robustness to noise, are a promising avenues for future studies.
We believe that our efforts demonstrate the promise of diffusion models for highly accurate reconstructions of physical flow fields given sparse measurements, an important open problem in different fields of engineering. Our work also highlights the importance of non-conflicting algorithms when training generative models with multiple objectives and physics constraints.
Our approach generally enourages further research into physics-constrained diffusion models for addressing probabilistic flow-related problems.
@@ Acknowledgements.
The data and source code that support the findings of this study are openly available in GitHub at https://github.com/tum-pbs/sparse-reconstruction.
@@
The authors have no conflicts to disclouse.
@@ Appendix A Comparison between point-wise and Gaussian masking methods
Table 6: Comparison between nearest interpolation, and our proposed masking methods. The evaluated task is the reconstruction of a field from 1.5625% of the known values. The used model was trained with the $\text{ConFIG}_{u}$ method.
@@ Appendix B Training details
To obtain comparable results, all 2D models were trained for 1000 epochs and a batch size of 5. A learning rate of $10^{-4}$ was used, with an exponential scheduler with $\gamma=0.98$, and a minimum $lr=10^{-5}$. The training of the 3D models was slightly different. Due to having a smaller number of samples we trained during 2000 epochs to achieve convergence. Since the samples are much bigger we needed a bigger GPU (NVIDIA L40S with 50 GB) for training and a slightly different batch size (4). We still used the same variable learning rate. The noise schedule $\beta$ for the diffusion forward process was kept as linear between $10^{-4}$ and 0.02.
The 2D network is based on a UNet model Ronneberger, Fischer, and Brox (2015), which is a convolutional neural network architecture that features a symmetric encoder–decoder structure with skip connections to capture both global context and fine details. To further expand the receptive field of the network, we also use self-attentionVaswani (2017) in the bottleneck layer.
As backbone model for 3D fields, we use the P3D model Holzschuh et al. (2025), a recent hybrid CNN-Transformer architecture for 3D physical simulations that can be applied in patches or to the whole sample. It follows a typical hierarchical U-shaped structure with an encoder and a decoder, based on a combination of convolutional and transformer blocks, which again combine the ideas from Swin transformers Liang et al. (2021) and diffusion transformers (DiT) Peebles and Xie (2023).
Appendix C Ablation of DDIM reverse steps

### v2

@@ IV Conclusions
The central idea of masked diffusion can be applied to super-resolution problems Fukami and Taira (2025, 2024); Page (2025), as these tasks are equivalent to sparse data reconstruction where the available data has a particular Cartesian structure, instead of being randomly placed. Applications additionally extend to experimental fluid mechanics, where a sparsely distributed set of probes can provide measurements at a limited number of locations, and the objective is to reconstruct the entire flow field. In such setups, obtaining complete flow field information is fundamentally constrained by the limitations of measurement techniques. Deploying a large number of probes is often infeasible, as intrusive sensors such as hot-wire anemometers, Pitot tubes, or pressure taps can disturb the flow and bias the measurements. Even advanced non-intrusive methods, such as Particle Image Velocimetry (PIV), are subject to trade-offs between field of view, resolution, and optical accessibility, resulting in data that is inherently sparse and noisy. In comparison to supervised models Guastoni et al. (2025), the use of a generative approach can improve the reconstruction of the small scales of the flow field.
We have shown that our method outperforms previous diffusion-based reconstruction approaches in both 2D and 3D scenarios. In particular, it excels at producing physically consistent reconstructions while preserving key flow statistics and accurately capturing the smallest scales. Despite that, it is subject to certain limitations inherent to any reconstruction approach, such as the need for specific data to train the underlying network. Moreover, in scenarios where frequent retraining of the model is necessary, our training method $\text{ConFIG}_{u}$ may not be ideal, as it is computationally more expensive than standard diffusion training. On the other hand, our approach outperforms previous diffusion-based works at inference time, as fewer reverse steps are needed.
Our results point to several interesting avenues for future work, such as evaluating how the proposed generative method performs for reconstructing other flow problems. We believe that our efforts demonstrate the promise of diffusion models for highly accurate reconstructions of physical flow fields given sparse measurements, an important open problem in different fields of engineering. Our work also highlights the importance of non-conflicting algorithms when training generative models with multiple objectives and physics constraints. Our approach generally encourages further research into physics-constrained diffusion models for addressing probabilistic flow-related problems.
@@ Acknowledgements.
The data and source code supporting the findings of this study are openly available on GitHub at https://github.com/tum-pbs/super-resolution.
@@
The authors have no conflicts to disclose.
@@ Appendix A Comparison between point-wise and Gaussian masking methods
Table 8: Comparison between nearest interpolation and our proposed masking methods. The evaluated task is the reconstruction of a field from 1.5625% of the known values. The used model was trained with the $\text{ConFIG}_{u}$ method.
@@ Appendix B Training details
The 2D network is based on a UNet model Ronneberger et al. (2015), which is a convolutional neural network architecture that features a symmetric encoder–decoder structure with skip connections to capture both global context and fine details. To further expand the receptive field of the network, we also utilize self-attention Vaswani (2017) in the bottleneck layer.
As a backbone model for 3D fields, we use the P3D model Holzschuh et al. (2025), a recent hybrid CNN-Transformer architecture for 3D physical simulations that can be applied in patches or to the whole sample. It follows a typical hierarchical U-shaped structure, comprising an encoder and a decoder, based on a combination of convolutional and transformer blocks. These blocks combine ideas from Swin Transformers Liang et al. (2021) and diffusion transformers (DiT) Peebles and Xie (2023).
To obtain comparable results, all 2D models were trained for 1000 epochs and a batch size of 5. A learning rate of $10^{-4}$ was used, with an exponential scheduler with $\gamma=0.98$, and a minimum $lr=10^{-5}$. The training of the 3D models was slightly different. Due to having a smaller number of samples, we trained for 2000 epochs to achieve convergence. Since the samples are much larger, we needed a more powerful GPU (NVIDIA L40S with 50 GB) for training and a slightly different batch size (4). We still used the same variable learning rate. The noise schedule $\beta$ for the diffusion forward process was kept as linear between $10^{-4}$ and 0.02.
Appendix C Ablation of interpolation methods for guidance in masked diffusion
Our masked diffusion method utilizes an interpolation of the sparse data to guide the reverse process. For simplicity, our first choice was to use nearest interpolation for this purpose. Nonetheless, one could use more advanced interpolation techniques. For example, we also considered linear interpolation and present a comparison on the 3D dataset in Tables 9 and 10. We observe slightly better results in terms of RMSE and residual error for linear interpolation, but a worse energy spectrum error. Additionally, computing a more advanced interpolation also requires extra runtime, which could be even higher than the runtime of the diffusion sampler itself. Therefore, we generally prefer to keep the nearest interpolation.
Models RMSE P-RMSE $\boldsymbol{L_{\text{{res}}}}$ LSiM E.spectrum
Masked diffusion ”nearest” (ours, default) 0.17 $\pm$ 0.01 0.11 $\pm$ 0.01 0.61 $\pm$ 0.03 0.06 $\pm$ 0.01 0.05 $\pm$ 0.01
Masked diffusion ”linear” (ours) 0.16 $\pm$ 0.01 0.20 $\pm$ 0.01 0.50 $\pm$ 0.02 0.06 $\pm$ 0.01 0.16 $\pm$ 0.01
Models RMSE P-RMSE $\boldsymbol{L_{\text{{res}}}}$ LSiM E.spectrum
Masked diffusion ”nearest” (ours, default) 0.20 $\pm$ 0.01 0.18 $\pm$ 0.01 0.66 $\pm$ 0.03 0.09 $\pm$ 0.01 0.09 $\pm$ 0.01
Masked diffusion ”linear” (ours) 0.19 $\pm$ 0.01 0.21 $\pm$ 0.01 0.53 $\pm$ 0.02 0.09 $\pm$ 0.01 0.18 $\pm$ 0.01
Appendix D Very sparse reconstruction with 0.1% input data
To support the results in Table 6 and Figure 15 with qualitative examples, we include visual reconstruction results in Figure 22. These visualizations show that the masked diffusion method provides plausible reconstructions despite the extremely sparse inputs, thanks to its ability to recreate the smallest structures in the fluid.
Figure 22: Example 3D isotropic turbulence 0.1% sparse data reconstruction (2D slices).
Appendix E Ablation of DDIM reverse steps